AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Wed May 10 19:06:18 2023
Author
*
:
Subject
*
:
<p>To test the detectors, adaptors were installed on the feed through flanges. A negative bias voltage was applied in steps of 10 V from 0 to 150 V. The V-I plots are attached for the four detectors. Three detectors - top right, top left, and bottom left (in the direction of the beam) - are fine. For the bottom right detector (in the direction of the beam), the current was fluctuating and was considerably high (1st test). The adaptors connecting to this detector were removed and checked by applying the bias voltage which showed zero current for voltages ranging from 0 to - 150 V, confirming the adaptors to be fine. They were then exchanged by other adaptors and the V-I test was repeated again (2nd test) with fluctuating and high current values. One of the flanges connecting to this detector was found to be light-tight. These tests were done at an ambient temperature of 21 degrees. Also, we couldn't find the rest of the three SHV to 2x Lemo 00 cables for the detector HV.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
detector_1.png
Original size: 1743x971
Attachment 2:
detector_2.png
Original size: 1743x971
Attachment 3:
detector_3.png
Original size: 1743x949
Attachment 4:
detector_4.png
Original size: 1743x996
Attachment 5:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown